# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu3,11*, Wenhao Huang2,3

### v2

@@ Appendix C Pre-training Details
We enable Flash Attention [99, 100] during pre-training. Gradient clipping is set to 1.0 for all stages. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW [101].
@@
Appendix D Supervised Fine-Tuning Details
@@
Language Only OpenHermes [102]
Multimodal ICL MMICL [103]
Multimodal CoT ScienceQA [104]
Chart Understanding Geo170K [105]
Instructional Image Generation InstructPix2Pix [106], MagicBrush [107]
ASR LibriSpeech [67], GigaSpeech [108], Common Voice [109]
Video Dialogue VideoChat2-IT [110]
Image QA Vision-Flan [111], VizWiz [59], LAION-GPT4V [^5] https://huggingface.co/datasets/laion/gpt4v-dataset, LLaVAR [112], OCR-VQA [113], VQA [57], TextVQA [114], OK-VQA [58], Mantis-Instruct [115]
Speech Generation SpeechInstruct [116]
Speech Understanding SpeechInstruct [116]
Image Captioning Flickr30K [62], MS-COCO [55]
Descriptive Image Generation Flickr30K [62], MS-COCO [55]
TTS GigaSpeech [108], Common Voice [109]
Video Generation MSR-VTT [117], MSVD [118]
Video Understanding MSR-VTT [117], MSVD [118], MSVD-QA [69], MSRVTT-QA [70]
Visual Storytelling VIST [50]
@@
The chat template is the same as used in Yi [47]. The system prompt is unified as: “You are MIO, an AI assistant capable of understanding and generating images, text, videos, and speech, selecting the appropriate modality according to the context.” except for speech generation and TTS whose system prompts are “You are MIO, an AI assistant capable of understanding images, text, videos, and speech, and generating speech. Please respond to the user with speech only, starting with <spch> and ending with </spch>.” to avoid randomness of the output modality.
@@
Similar to pre-training (c.f., Appendix C), we enable Flash Attention [99, 100] during supervised fine-tuning. Gradient clipping is set to 1.0. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW [101].
Appendix E Evaluation Details.
@@
Image QA (We use the prompt templates in LMMs-Eval [119]).
@@
Appendix F More Demonstrations.
Figure 4: Multimodal Chain-of-Thought and Multimodal In-Context Learning Demos.

### v3

@@ Appendix B Pre-training Details
We enable Flash Attention (Dao et al., 2022; Dao, 2023) during pre-training. Gradient clipping is set to 1.0 for all stages. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov & Hutter, 2017).
@@
Appendix C Supervised Fine-Tuning Details
@@
Language Only OpenHermes (Teknium, 2023)
Multimodal ICL MMICL (Zhao et al., 2023)
Multimodal CoT ScienceQA (Lu et al., 2022)
Chart Understanding Geo170K (Gao et al., 2023)
Instructional Image Generation InstructPix2Pix (Brooks et al., 2023), MagicBrush (Zhang et al., 2024)
ASR LibriSpeech (Panayotov et al., 2015), GigaSpeech (Chen et al., 2021), Common Voice (Ardila et al., 2020)
Video Dialogue VideoChat2-IT (Li et al., 2023c)
Image QA Vision-Flan (Xu et al., 2023), VizWiz (Gurari et al., 2018), LAION-GPT4V [^6] https://huggingface.co/datasets/laion/gpt4v-dataset, LLaVAR (Zhang et al., 2023c), OCR-VQA (Mishra et al., 2019), VQA (Goyal et al., 2016), TextVQA (Singh et al., 2019), OK-VQA (Marino et al., 2019), Mantis-Instruct (Jiang et al., 2024)
Speech Generation SpeechInstruct (Zhang et al., 2023a)
Speech Understanding SpeechInstruct (Zhang et al., 2023a)
Image Captioning Flickr30K (Plummer et al., 2015), MS-COCO (Lin et al., 2014)
Descriptive Image Generation Flickr30K (Plummer et al., 2015), MS-COCO (Lin et al., 2014)
TTS GigaSpeech (Chen et al., 2021), Common Voice (Ardila et al., 2020)
Video Generation MSR-VTT (Xu et al., 2016), MSVD (Chen & Dolan, 2011b)
Video Understanding MSR-VTT (Xu et al., 2016), MSVD (Chen & Dolan, 2011b), MSVD-QA (Chen & Dolan, 2011a), MSRVTT-QA (Xu et al., 2017)
Visual Storytelling VIST (Huang et al., 2016)
@@
The chat template is the same as used in Yi (AI et al., 2024). The system prompt is unified as: “You are MIO, an AI assistant capable of understanding and generating images, text, videos, and speech, selecting the appropriate modality according to the context.” except for speech generation and TTS whose system prompts are “You are MIO, an AI assistant capable of understanding images, text, videos, and speech, and generating speech. Please respond to the user with speech only, starting with <spch> and ending with </spch>.” to avoid randomness of the output modality.
@@
Similar to pre-training (c.f., Appendix B), we enable Flash Attention (Dao et al., 2022; Dao, 2023) during supervised fine-tuning. Gradient clipping is set to 1.0. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov & Hutter, 2017).
Appendix D Evaluation Details.
@@
Image QA (We use the prompt templates in LMMs-Eval (Li* et al., 2024)).
@@
Appendix E More Experiments
E.1 Image Generation Evaluation
We compute two additional automatic metrics for evaluating image generation, i.e., SSIM (Wang et al., 2004) and Aesthetic Predictor v2.5777https://github.com/discus0434/aesthetic-predictor-v2-5?tab=readme-ov-file for the evaluation of structural integrity and aesthetics, respectively. SSIM (Structural Similarity Index Measure) evaluates the perceptual similarity between the generated images and the ground-truth images, focusing on luminance, contrast, and structure, with scores ranging from -1 (dissimilar) to 1 (identical). Aesthetic Predictor V2.5 is a SigLIP (Zhai et al., 2023)-based predictor that evaluates the aesthetics of an image on a scale from 1 to 10 (10 is the best). In addition, we randomly select 100 image descriptions from MS-COCO test set, and used each model to generate images accordingly for human preference evaluation. We ask 3 annotators to rank 3 images generated by the 3 models: “given the image description, which image is preferred?” The average ranking of MIO’s, AnyGPT’s, and Emu’s generated images are 1.2 (MIO), 2.9 (AnyGPT), 1.9 (Emu). MIO aligns the best with the human preference. The percentage agreement between the three annotators (calculated as the number of cases with identical rankings by all annotators divided by 100) is 82.3%, indicating a high consistency in the human evaluation.
Dataset MS-COCO Flickr30K MS-COCO Subset
Metric SSIM ($\uparrow$) Aesthetic ($\uparrow$) SSIM ($\uparrow$) Aesthetic ($\uparrow$) Human Avg. Ranking ($\downarrow$)
Emu 0.1749 3.733 0.1451 3.893 1.9
AnyGPT 0.1960 3.954 0.1585 4.251 2.9
MIO 0.2307 4.019 0.1727 4.326 1.2
Table 12: Image generation evaluation by SSIM, Aesthetic Predictor V2.5, and human preference.
E.2 Speech-to-Speech Evaluation
Model Supported Workflow Content Score (1-5 points) ($\uparrow$)
MIO s2s 1.4
[…]
(Fang et al., 2024)
s2t→t2s 2.4
AnyGPT s2t→t2s 1.8
Table 13: Speech-to-Speech performance. “s2s” means “speech-to-speech”, while “s2t” and “t2s” denote “speech-to-text” and “text-to-speech”, respectively.
Since there is a lack of speech to speech evaluation benchmarks, we randomly sample some conversations from the moss-002-sft dataset [^8] https://huggingface.co/datasets/fnlp/moss-002-sft-data and convert them into speech-to-speech format. Following the evaluation procedures outlined in LLaMA-Omni (Fang et al., 2024), we use the content score metric obtained from GPT-4o (OpenAI et al., 2024) to assess whether the model’s response effectively addresses the user’s instructions. The results are shown in Table 13.
Though the content score of MIO is slightly lower than LLaMA-Omni and AnyGPT, both LLaMA-Omni and AnyGPT first generate text replies and then convert these into voice. However, our model, MIO, is capable of directly generating speech responses to speech queries.
E.3 TTS Evaluation
Model GLOBE LibriSpeech test-clean
WER ($\downarrow$) Speech Similarity ($\uparrow$) WER ($\downarrow$) Speech Similarity ($\uparrow$)
MIO 9.8 67.8 10.3 75.1
AnyGPT 27.9 67.3 28.1 71.3
Table 14: More automatic evaluations for the TTS performance.
We select two additional benchmarks, LibriSpeech test-clean (Panayotov et al., 2015) and GLOBE (Wang et al., 2024b), to evaluate the performance of TTS between our model and AnyGPT. For fair comparison, we don’t specify the input voice prompt during evaluation of MIO and AnyGPT. WER (Word Error Rate) and speaker similarity are employed as the automatic metrics. The results are shown in Table 14. The results show that MIO performs significantly better than AnyGPT on both WER and speaker similarity across both benchmarks.
Table 15: Human evaluation for the TTS performance.
MIO Win 54%
Tie 25%
MIO Lose 21%
Additionally, we conduct a human evaluation to assess the speech quality of the outputs from MIO and AnyGPT. In this evaluation, participants are provided with the target speech, the speech generated by AnyGPT, and the speech generated by our model. They are tasked with determining which one sounded more natural and closer to the target speech. Evaluators could choose one of the two generated speeches or indicate that they find them equally natural. Each evaluation is rated by three independent human evaluators, and we report the average scores. The results are shown in Table 15. MIO significantly outperforms AnyGPT in the human evaluation, consistent with the results from the automatic evaluation.
E.4 Loss Curves
Figure 3: Loss curves of pretraing stages.
We plot the loss curves for each stage in Figure 3. We can observe that when introducing a new data type (i.e., image-text interleaved data) in stage 2, the training loss suddenly increases. However, in the third pretraining stage, i.e., the speech-enhancement stage, the training loss transitions more smoothly. Despite the fluctuations in loss between stages, which do have some impact on downstream performance during the fluctuation periods, we find that with continued training, the model’s loss quickly recovers to its previous convergence level and continues optimizing effectively.
E.5 More Demonstrations.
Figure 4: Demonstrations of MIO’s advanced abilities. Yellow: inputs; Green: outputs.
Figure 5: Demonstrations of MIO’s basic abilities. Yellow: inputs; Green: outputs.
Figure 6: Multimodal Chain-of-Thought and Multimodal In-Context Learning Demos. Yellow: inputs; Green: outputs.